OpenStudy (anonymous):

PLZ HELP! What would the equation that represents the line passing through the points (1, 4) and (-2, 13) look like?

OpenStudy (anonymous):

y=mx+b lets use the two coordinates to find the values of m and b: (1,4) and (-2,13) m=(y2-y1)/(x2-x1)=(13-4)/(-2-3)=9/-3=-3 b: y=mx+b---> 4=(-3)(1)+b--->b=4+3=7 equation: y=-3x+7